### v2

# The trade-off between diagonal and off-diagonal elements in the eigenstate thermalization hypothesis

Abstract
To bypass using local observables as intermediate quantities in proving the eigenstate thermalization hypothesis (ETH), we have introduced an observable-independent measure of distinguishability. In this paper, we establish the connection between this measure and several other ETH measures in a more natural way. We first demonstrate a universal trade-off relation between the diagonal and off-diagonal elements of the measure. We then extend this discussion to eigenstate typicality and the average observable. This trade-off relationship reveals that the exponential growth of off-diagonal elements directly suppresses their own values, as well as indirectly suppressing the diagonal elements. This provides a new perspective on the physical mechanisms underlying ETH. Finally, through numerical calculations on a one-dimensional Ising spin chain, we explore various trade-off relationships and examine strong and weak ETH.
I Introduction
The equilibration and thermalization of an isolated quantum system are crucial for understanding how quantum statistical mechanics emerges from unitary quantum mechanics. The eigenstate thermalization hypothesis (ETH) [1, 2] posits that the expectation values of quantum observables in an energy eigenstate should approximately match the thermal expectation values, offering a potential mechanism for the thermalization of such systems. In the original proposal by Deutsch and Srednicki [3, 4, 5], a random perturbation is introduced to a closed quantum system, and the ETH is expected to hold if the perturbed system becomes chaotic. However, this randomness complicates the understanding of its physical origins and mathematical description.
In many studies of the ETH, such as [8, 9, 6, 10, 11, 14, 12, 13, 7], one does not assume random energy perturbations or random Hamiltonians but instead aim to derive statistical properties solely from quantum properties. These studies focus on demonstrating the so-called ”strong (weak) ETH”. The strong (weak) ETH states that all (or almost all) energy eigenstates exhibit thermal properties with respect to local observables, and it postulates that the local distinguishability between an eigenstate and a universal density matrix decreases exponentially with system size. It is generally believed that strong ETH does not apply to integrable or many-body localized systems. On the other hand, the weak ETH emphasizes eigenstate typicality, suggesting that the average distinguishability between an eigenstate and the canonical ensemble decreases algebraically with system size, and it can be proven for generic translationally invariant systems, including integrable systems [9, 11, 12].
In many mathematical proofs of the ETH, such as [15, 11, 12], they typically define a distinguishability measure based on an arbitrary observable and then demonstrate that the distinguishability between the eigenstate and the canonical ensemble approaches zero in the thermodynamic limit, regardless of the observable’s specific form. This approach motivates us to develop a distinguishability measure independent of observables. The trace norm used in [14] serves as such a measure, but it is challenging to transform and less convenient for further theoretical analysis. In our previous work [16], we explored ETH using Belavkin-Staszewski relative entropy and the variance $V(\rho_{B},O^{B_{1}}_{ij})$. However, we did not fully appreciate the significance of certain formulas or conduct numerical calculations on specific models, leading to conclusions similar to the weak ETH. In this paper, we further investigate the properties of $V(\rho_{B},O^{B_{1}}_{ij})$ and conduct numerical verification with specific models.
We begin in Section II with preliminary results about ETH, subsystem ETH, weak ETH and ETH with the average observable. We show how the distinguishability measure $V(\rho_{B},O^{B_{1}}_{ij})$ relates to commonly used measures and demonstrate how it can be extended to the average observable. In Section III, we prove the existence of a universal trade-off equation for the diagonal and off-diagonal elements of $V(\rho_{B},O^{B_{1}}_{ij})$. This trade off relationship can be extended to the discussion of eigenstate typicality and the average observable. In Section IV, we examine the trade-off relation and the strong (weak) ETH by numerically simulating an Ising spin chain with a transverse and longitudinal magnetic field. Finally, in Section V, we conclude the paper and discuss related issues.
II Preliminaries
In this section, we review the basics of ETH and subsystem ETH, including weak ETH with eigenstate typicality as discussed in [6, 10, 11, 7], and eigenstate thermalization with macroscopic observables as explored in [11, 12].
II.1 ETH and subsystem ETH
Consider an isolated or closed quantum system $B$ with Hamiltonian $H$. Suppose the system has size $N$ and $H$ has eigenvectors |$\ket{E_{i}},i=1,2,...,D_{N}$ with energy eigenvalues $E_{i}$, such that $H\ket{E_{i}}=E_{i}\ket{E_{i}}$. For a few-body observable $A^{B_{1}}$, the measure
$d^{(1)}(\Pi^{ij}_{B},\rho_{B};A^{B_{1}}):=\lvert\text{Tr}[A^{B_{1}}(\ket{E_{i}}\bra{E_{j}}-\rho_{B}\delta_{ij})]\rvert$ | (1)
quantifies the local (in)distinguishability between the energy eigenstates and the density matrix $\rho_{B}$. According to local ETH [1], this measure is expected to satisfy $d^{(1)}(\Pi^{ij}_{B},\rho_{B}(E);A^{B_{1}})\sim\mathcal{O}(e^{-S(E)/2})$, where $\rho_{B}(E)$ is some universal density matrix that changes slowly with energy $E=(E_{i}+E_{j})/2$.
In the measure (1), the distinguishability depends on the chosen observable. For an observable-independent measure, a more straightforward approach is to consider the following measure:
$d^{(2)}(\sigma^{B_{1}}_{ij},\rho_{B_{1}}):=\lVert\sigma^{B_{1}}_{ij}-\rho_{B_{1}}\delta_{ij}\rVert,$ (2)
where $\sigma^{B_{1}}_{ij}=\text{Tr}_{\bar{B}_{1}}\ket{E_{i}}\bra{E_{j}}$ | is the reduced matrix on the subsystem $B_{1}$ and the trace distance $\lVert O\rVert=\frac{1}{2}\text{Tr}\sqrt{O^{\dagger}O}$. The measure (2) is used to study subsystem ETH, which posits that $d^{(2)}(\sigma^{B_{1}}_{ij},\rho_{B_{1}}(E))\sim\mathcal{O}(e^{-S(E)/2})$. Subsystem ETH, as defined by (2), is actually stronger than local ETH as described in (1). This is supported by the following inequality [14]
$d^{(1)}(\Pi^{i}_{B},\rho_{B};A_{B_{1}})\leqslant\sqrt{\lVert\sigma^{B_{1}}_{i}-\rho_{B_{1}}\rVert_{1}\text{Tr}[(\Pi^{i}_{B}+\rho_{B})A_{B_{1}}^{2}]},$ (3)
where $\Pi^{i}_{B}=\ket{E_{i}}\bra{E_{i}}$ |, $\sigma_{i}=\sigma_{ii}$ and $\lVert\cdot\rVert_{1}=2\lVert\cdot\rVert$ is the Schatten $1$-norm (trace norm).
Although straightforward, the measure (2) is not in a very convenient form and is challenging to relate to other quantities. Using the Schatten $2$-norm instead of the trace norm yields a weaker condition, making it difficult to ensure that $d^{(1)}$ remains small. We introduce an alternative, less direct measure. By applying the rescaling map $\mathcal{J}^{\alpha}_{\rho}(\cdot):=\rho^{\alpha}(\cdot)\rho^{\alpha}$, we transform the elements in eq. 2 as follows:
$\sigma^{B_{1}}_{ij}-\rho_{B_{1}}\delta_{ij}=\mathcal{J}_{\rho_{B_{1}}}^{1/2}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}}),$ (4)
where $O^{B_{1}}_{ij}:=\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\sigma^{B_{1}}_{ij})$ and $\braket{O^{B_{1}}_{ij}}=\text{Tr}(O^{B_{1}}_{ij}\rho_{B})=\delta_{ij}$. Note that the rescaling map involves the inverse $\rho_{B_{1}}^{-1}$, requiring that the density matrix $\rho_{B_{1}}$ be strictly positive, which holds for the canonical and microcanonical ensembles considered here. Applying Hölder’s inequality, we obtain
$2d^{(2)}(\sigma^{B_{1}}_{ij},\rho_{B_{1}})=\lVert\mathcal{J}_{\rho_{B_{1}}}^{1/2}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}})\rVert_{1}$
$\leq\lVert(\rho_{B_{1}})^{1/2}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}})\rVert_{2}\lVert(\rho_{B_{1}})^{1/2}\rVert_{2}$
$=\sqrt{\text{Tr}[(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}})^{\dagger}\rho_{B}(O^{B_{1}}_{ij}-\braket{O^{B_{1}}_{ij}})]}$
$=\sqrt{V(\rho_{B},O^{B_{1}}_{ij})}=:d^{(3)}(\sigma^{B_{1}}_{ij},\rho_{B_{1}}),$ (5)
where $\lVert T\rVert_{p}=\{\text{Tr}[(T^{\dagger}T)^{p/2}]\}^{1/p}$ denotes the Schatten $p$-norm and the variance is defined as:
$V(\rho,A):=\text{Tr}[(A-\braket{A})^{\dagger}\rho(A-\braket{A})].$ (6)
In this paper, we will focus on the measure $d^{(3)}$. As indicated by section II.1, $d^{(3)}$ serves as an upper bound for the other measures, making it the strongest ETH measure. Additionally, because variance is easily transformed, it facilitates connections with other quantities, which will be demonstrated later.
II.2 Weak ETH with eigenstate typicality
It is a strong assumption that all energy eigenstates satisfy the ETH. People are usually concerned with the Hilbert subspace within an ensemble. Consider the state of the total system $B$, denoted by $\rho_{B}=\sum_{j}p_{j}\Pi^{j}_{B}$. In many investigations of the “weak” ETH, the goal is to derive statistical properties and quantify probabilistic typicality or concentration with respect to the measure $d^{(1)}$. For this, the following expression is used [16]
$\Delta(\rho_{B},A)=\int[d^{(1)}(\Pi^{i}_{B},\rho_{B};A^{B_{1}})]^{2}p_{d},$ (7)
where the probability distribution $p_{d}$ is derived from the $p_{j}$’s through a Radon-Nikodym derivative. According to the Chebyshev inequality, we have
$P_{\rho}(d^{(1)}(\Pi^{i}_{B},\rho_{B};A^{B_{1}})\geqslant\epsilon)\leqslant\frac{\Delta(\rho,A)^{2}}{\epsilon^{2}},\quad\forall\epsilon\in\mathbb{R^{+}}.$ (8)
Thus, when $\Delta(\rho,A)$ is very small, the expectation of the projectively measured observable would concentrate on the expectation of observable calculated with respect to the state $\rho_{B}$. If
$\Delta(\rho^{\text{mc}},A_{B_{1}})\sim\mathcal{O}(N^{-\alpha}),$ (9)
with $10<\alpha<$ 1, meaning that the expectations of a local observable $A^{B_{1}}$ with respect to the results of local measurements concentrate around the expectation of $A^{B_{1}}$ with respect to $\rho^{\text{mc}}$, then it follows that each pure state $\Pi^{i}_{B}$ cannot be distinguished from the microcanonical state $\rho^{\text{mc}}$ in the large $N$ limit. This result characterizes the weak ETH with eigenstate typicality [6, 11]. By substituting the measure $d^{(1)}$ with other measures, different forms of the weak ETH can be obtained.
II.3 The average observable
In a many-body system, the states of individual subsystems exhibit fluctuations. Even in a system with a Hamiltonian that is translationally invariant, only certain states, such as the Gibbs state (canonical ensemble) and the microcanonical ensemble, maintain guaranteed translational invariance. The energy eigenstates themselves can still fluctuate. Consequently, compared to the translationally invariant ensemble, spatial fluctuations can make the energy eigenstates more distinguishable.
Consider partitioning the lattice of system $B$ into $\mathcal{C}=\lvert B\rvert/\lvert B_{1}\rvert$ blocks of equal size, where $\lvert B\rvert$ denotes the number of lattice points in $B$. These $\mathcal{C}$ blocks are identical copies of $B_{1}$. In order to eliminate the influence of spatial fluctuations, or consider the macroscopic observable [12], one can use the average observable
$A^{B}=\frac{1}{\mathcal{C}}\sum_{i}A^{B_{i}},$ (10)
where $A^{B_{i}}$ represents the translational copy of $A^{B_{1}}$. Substituting $A^{B}$ into eq. 1 gives
$d^{(1)}(\Pi^{ij}_{B},\rho;A^{B})=\lvert\frac{1}{\mathcal{C}}\sum_{k}\text{Tr}[(\sigma^{ij}_{B_{k}}-\delta_{ij}\rho_{B_{k}})A^{B_{k}}]\rvert$
$=\lvert\text{Tr}[(\frac{1}{\mathcal{C}}\sum_{k}\sigma^{ij,k}_{B_{1}}-\delta_{ij}\overline{\rho}_{B_{1}})A^{B_{1}}]\rvert$
$=d^{(1)}(\frac{1}{\mathcal{C}}\sum_{k}\sigma^{ij,k}_{B_{1}},\overline{\rho}_{B_{1}};A^{B_{1}}),$ (11)
where $\sigma^{j,k}_{B_{1}}$ are the translational copies of $\sigma^{j}_{B_{k}}$, $\overline{\rho}_{B}=(\sum_{k}\hat{T}_{k}\rho_{B})/\mathcal{C}$ and $\hat{T}_{k}$ is the translation operator that maps $B_{k}$ to $B_{1}$. According to our previous work [16], an upper bound can be given by the variance
$d^{(1)}(\Pi^{ij}_{B},\rho;A)^{2}\leqslant V(\overline{\rho},\overline{O}^{B_{1}}_{ij})\lVert A^{B_{1}}\rVert_{\infty}^{2},$ (12)
where $\overline{O}^{B_{1}}_{ij}=\frac{1}{\mathcal{C}}\sum_{k}\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\sigma^{ij,k}_{B_{1}})$, and $\sigma^{ij,k}_{B_{1}}$ are the translational copies of $\sigma^{ij}_{B_{k}}$.
III Trade-off between diagonal and off-diagonal elements
III.1 Trade-off between diagonal and off-diagonal elements
The diagonal element $V_{\text{dg}}^{i}=V(\rho_{B},O^{B_{1}}_{ii})$ and the off-diagonal element $V_{\text{off}}^{ij}=V(\rho_{B},O^{B_{1}}_{ij;i\neq j})$ in $d^{(3)}$ are not independent. According to appendix A, they are related by the following equation
$V_{\text{dg}}^{i}+\sum_{j;j\neq i}V_{\text{off}}^{ij}=\text{Tr}(\rho_{B_{1}}^{-1})-1,$ (13)
where the index $j$ sums over the entire Hilbert space except $i$, and its dimension $D_{N}-1$ grows exponentially with the number of sites. In contrast, the right-hand side of (13) depends solely on the local state of $\rho_{B}$, generally increasing with $D_{B_{1}}^{2}$ but remaining relatively constant with $N$. This causes the off-diagonal elements to be exponentially suppressed. Considering the average of these off-diagonal elements
$V_{\text{off}}^{i}=\frac{1}{D_{N}-1}\sum_{j;j\neq i}V_{\text{off}}^{ij},$ (14)
which will decrease exponentially with $N$. However, this exponential decrease in the average of the off-diagonal elements does not imply that the maximum values of the off-diagonal elements and the diagonal element will also decrease exponentially. Although the diagonal elements are not directly suppressed, eq. 13 provides a trade-off relation between the diagonal element and the off-diagonal elements. If $V_{\text{dg}}^{i}/V_{\text{off}}^{i}=f^{i}(N)$, then we have
$V_{\text{dg}}^{i}=\frac{f^{i}(N)}{f^{i}(N)+D_{N}-1}(\text{Tr}(\rho_{B_{1}}^{-1})-1).$ (15)
On one hand, if $f^{i}(N)$ does not increase exponentially, or grows exponentially at a rate much slower than $D_{N}$, then in the large $N$ limit, the diagonal elements will also be exponentially small. On the other hand, if the diagonal element is not decreasing exponentially, $f^{i}(N)$ must be growing exponentially. Thus, the exponential growth of $f^{i}(N)$ is a necessary condition for the diagonal element to avoid exponential decay.
In this paper, we do not delve further into how interactions affect the growth of $f^{i}(N)$ with $N$. However, as will be evident from the subsequent numerical calculations, the influence of integrability on this growth is crucial. To fully address the underlying theoretical mechanisms and other factors, such as the choice of $\rho_{B_{1}}$, further investigation is needed. One potential approach involves using Hölder’s inequality
$V_{\text{dg}}^{i}=\lVert(\sigma^{i}_{B_{1}}-\rho_{B_{1}})\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}$
$\leq\lVert(\sigma^{i}_{B_{1}}-\rho_{B_{1}})(\sigma^{ij}_{B_{1}})^{-1}\rVert^{2}_{\infty}\lVert\sigma^{ij}_{B_{1}}\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}.$ (16)
Combining this result with eq. 14, it can be easily shown that
$f^{i}(N)\leq\frac{D_{N}-1}{\sum_{j;j\neq i}\lVert(\sigma^{i}_{B_{1}}-\rho_{B_{1}})(\sigma^{ij}_{B_{1}})^{-1}\rVert^{-2}_{\infty}}.$ (17)
This inequality’s right-hand side may provide a more straightforward framework to discuss the effects of interactions and the choice of the $\rho_{B_{1}}$ on $f^{i}(N)$.
Figure 1: Diagonal and off-diagonal elements versus eigenstate. The eigenstate indices are arranged from low to high energy. In the plot, $\text{Var}_{\text{dg}}=V_{\text{dg}}^{i}$, $\text{Var}_{\text{off}}=\sum_{j;j\neq i}V_{\text{off}}^{ij}$, $\text{SN}_{\text{dg}}=\lVert\sigma^{B_{1}}_{i}-\rho_{B_{1}}\rVert_{1}^{2}$ and $\text{SN}_{\text{off}}=\sum_{j;j\neq i}\lVert\sigma^{B_{1}}_{ij}\rVert_{1}^{2}$. The measure $\text{Var}_{\text{dg}}$ closely aligns with $\text{SN}_{\text{dg}}$, making the data points for $\text{Var}_{\text{dg}}$ overlap with those for $\text{SN}_{\text{dg}}$. Parameters: $g=1.05$, $h=0.1$, $N=12$, $N_{B_{1}}$=1, $\rho_{B}=\sum_{i}\Pi_{i}/D_{N}$.
The index $i$ in eq. 13 can be chosen arbitrarily, but the right side of the formula is independent of $i$. When the measure $V_{\text{dg}}$ of a given eigenstate is large, the average measure of its off-diagonal elements $V_{\text{off}}^{i}$ must be small, and vice versa. This trade-off is illustrated in fig. 1. Since the measure $d^{(2)}$ and $d^{(3)}$ are related by section II.1, the trade-off between the diagonal and off-diagonal elements is also reflected in $d^{(2)}$, although it is not an exact equation relationship like eq. 13. Calculations using a specific model are shown in fig. 1, where the diagonal element and the sum of the corresponding off-diagonal elements are compared. From the fig. 1, we observe that the diagonal element is generally much larger than the average of the off-diagonal elements. However, because there are $D_{N}-1$ off-diagonal elements, their total sum significantly exceeds that of the diagonal elements.
The formula (13) holds for any $\rho_{B}$ as long as $\rho_{B_{1}}$ is strictly positive. This includes not only the canonical ensemble $\rho^{\text{c}}=e^{-\beta H}/Z(\beta)$ and microcanonical ensembles, but also pure states such as $\Pi_{i}^{B}=\ket{E_{i}}\bra{E_{i}}$ |. In the case of pure states, the diagonal element in eq. 13 will be zero, leading to
$\sum_{j;j\neq i}V(\Pi_{i}^{B},O^{B_{1}}_{ij})=\text{Tr}(\sigma_{B_{1}}^{-1})-1,$ (18)
where the sum of all corresponding off-diagonal elements is directly determined by the local state $\sigma_{B_{1}}=\text{Tr}_{\bar{B}_{1}}\Pi_{i}^{B}$.
III.2 Weak ETH
We examine the use of measure $d^{(3)}$ to study the weak ETH with eigenstate typicality by considering both ”diagonal” and ”off-diagonal” probabilistic typicalities. The ”diagonal” probabilistic typicality is given by
$\braket{V_{\text{dg}}}:=\sum_{i}p_{i}V_{\text{dg}}^{i}$ (19)
while the ”off-diagonal” probabilistic typicality is
$\braket{V_{\text{off}}}:=\sum_{i}p_{i}V_{\text{off}}^{i}.$ (20)
Given that the relationship in eq. 13 holds for each $i$, averaging over the probabilities yields
$\braket{V_{\text{dg}}}+(D_{N}-1)\times\braket{V_{\text{off}}}=\text{Tr}(\rho_{B_{1}}^{-1})-1.$ (21)
As discussed previously, $\braket{V_{\text{off}}}$ decreases exponentially with $N$. The behavior of $\braket{V_{\text{dg}}}$ depends on the specific model.
III.3 The average observable
If we consider the distinguishability measure $V(\overline{\rho}_{B},\overline{O}^{B_{1}}_{ij})$ given by the average observable, then according to appendix A, the corresponding diagonal and off-diagonal elements exhibit the following trade-off relationship
$V(\rho_{B},\overline{O}^{B_{1}}_{ii})+\sum_{j;j\neq i}V(\rho_{B},\overline{O}^{B_{1}}_{ij})=\frac{1}{\mathcal{C}}(\text{Tr}(\rho_{B_{1}}^{-1})-1)$
$+\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,l,k;k\neq l}p^{\prime}_{\alpha}\text{Tr}[O^{B_{k},{\dagger}}_{\alpha\beta,1}\otimes O^{B_{l}}_{\alpha\beta,1}(\sigma^{i}_{B_{k}B_{l}}-\sigma^{i}_{B_{k}}\otimes\rho^{1}_{B_{l}})],$ (22)
where $\rho^{1}_{B_{l}}$ is the translational copy of $\rho_{B_{1}}$. Although formula (III.3) is valid for any $\rho_{B}$, it connects to eq. 12 specifically when using $\overline{\rho}_{B}$. Thus, all quantities related to $\rho_{B}$ should be replaced with $\overline{\rho}_{B}$, which is translation invariant. In this case $\overline{\rho}^{1}_{B_{l}}=\overline{\rho}_{B_{l}}$. The local term in section III.3 becomes $\text{Tr}(\overline{\rho}_{B_{1}}^{-1})-1$, which is related to the local state $\overline{\rho}_{B_{1}}$ and includes an additional factor of $1/\mathcal{C}$. The local term decreases algebraically as $\mathcal{C}$ increases. The second term on the right side of section III.3 is related to the spatial correlation of $\sigma^{i}$ and the difference between $\sigma^{i}_{B_{l}}$ and $\overline{\rho}_{B_{l}}$.
We can also consider both the average observable and eigenstate typicality. In this case, we need to examine the following “diagonal” probabilistic typicality
$\braket{V_{\text{dg}}}=\sum_{i}p_{i}V(\rho,\overline{O}^{B_{1}}_{ii}).$ (23)
And the “off-diagonal” probabilistic typicality
$\braket{V_{\text{off}}}=\frac{1}{D_{N}-1}\sum_{i,j;j\neq i}p_{i}V(\rho,\overline{O}^{B_{1}}_{ij}).$ (24)
After averaging the probability of section III.3, we obtain the following trade-off relationship
$\braket{V_{\text{dg}}}+(D_{N}-1)\times\braket{V_{\text{off}}}=\frac{1}{\mathcal{C}}(\text{Tr}(\rho_{B_{1}}^{-1})-1)+V^{\text{Avg}}_{\text{Cor}},$ (25)
where
$V^{\text{Avg}}_{\text{Cor}}:=\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,l,k;k\neq l}p^{\prime}_{\alpha}\text{Tr}[O^{B_{k},{\dagger}}_{\alpha\beta}\otimes O^{B_{l}}_{\alpha\beta}(\rho_{B_{k}B_{l}}-\rho_{B_{k}}\otimes\rho_{B_{l}})].$ (26)
In formula (24), except for $\rho_{B_{k}B_{l}}$ and $\rho_{B_{k}}$, all quantities related to $\rho_{B}$ need to be replaced with $\overline{\rho}_{B}$. If $\rho_{B}$ is translation invariant, then $\overline{\rho}_{B}=\rho_{B}$, and the second term on the right side of eq. 25 is solely related to the correlation of $\rho_{B}$. If the spatial correlation of $\rho_{B}$ decays exponentially or algebraically fast enough, the right side of eq. 25 will exhibit algebraic decay of the form $O(N^{-\epsilon})$ [16]. Compared to eq. 21, the right side of eq. 25 depends not only on the local state but also on the rate of correlation decay. Both terms generally decrease algebraically with increasing $N$, consistent with the intuition that spatial averaging smooths and reduces differences. Although this algebraic decay alone does not guarantee strong ETH, it supports the validity of weak ETH even for integrable systems. This algebraic decay aligns with the findings of our previous article [16], though we did not previously account for the $D_{N}-1$ off-diagonal elements in eq. 25, which cause the average off-diagonal elements to be exponentially suppressed and further restrict their values. Consequently, as the right side of eq. 25 decreases with $N$, the exponential decrease of $\braket{V_{\text{off}}}$ remains applicable. Additionally, due to the smoothing effect of the average observable, the values of $\braket{V_{\text{dg}}}$ and $\braket{V_{\text{off}}}$ here will be smaller than those defined in the previous section.
Figure 2: The correlation term $V^{\text{Avg}}_{\text{Cor}}$ versus $N$. Parameters: $g=1.05$, $N_{B_{1}}=1$.
IV Numerical Results
To illustrate the previous proof, we numerically simulate an Ising spin chain with both transverse and longitudinal magnetic fields, described by the Hamiltonian
$H=\sum_{k=1}^{N}(-\sigma_{z}^{k}\otimes\sigma_{z}^{k+1}+g\sigma^{k}_{x}+h\sigma_{z}^{k}).$ (27)
This system is nonintegrable unless one of the coupling constants $g$ or $h$ is zero. For simplicity, we apply periodic boundary conditions, although the translation-invariance of the system in the numerical simulation is not a requirement for our results. Almost all the derivations in sections II and III are applicable to non-translation-invariant systems as well.
We used ITensor Library [17] for exact diagonalization and numerical calculations of the system. To facilitate comparison with the results in [14], we consider the cases $g=1.05$ with $h=0$ or $h=0.1$. Basic properties of the system, including energy state distribution, energy-temperature curve, and mutual information of the canonical ensemble, are presented in appendix B.
IV.1 The impact of energy difference and temperature on distinguishability
Figure 3: Parameters: $g=1.05$, $N=15$, $N_{B_{1}}=1$. (a) The local diagonal element measure $V_{\text{dg}}^{i}$ versus energy (difference). We used $\rho_{B}=\sum_{i}\Pi_{i}/D_{N}$ and $E(\rho_{B})=0$. (b) The average $V_{\text{dg}}^{i}$ under the energy shell with different energy centers $E$. Here $\Delta E=0.2$ and $\rho_{B}=\rho^{\text{c}}(\beta)$ with $E(\beta)=E$.
Although equation (13) holds for all energy eigenstates, the measure $V_{\text{dg}}^{i}$ for the diagonal elements varies across different energy eigenstates. This difference is evident in fig. 1. As shown in fig. 3(a), there is some randomness in the measure $V_{\text{dg}}^{i}$ with respect to the eigenstate, but generally, the greater the energy difference between $E_{i}$ and $E(\rho_{B})$, the higher the average distinguishability. In addition, integrability also has a great influence. For the integrable systems, the diagonal elements $V_{\text{dg}}^{i}$ tend to be more variable, and the average distinguishability is typically higher.
Given these factors, it is not always meaningful to consider the maximum value of $V_{\text{dg}}^{i}$ over the entire energy eigenstate space for a fixed $\rho_{B}$. Instead, it is more insightful to examine the maximum value within a specific energy shell, such as
$\mathcal{M}_{E,\Delta}=(E-\Delta E/2,E+\Delta E/2),$ (28)
with $E=E(\rho_{B})$. Moreover, when analyzing energy eigenstates within $\mathcal{M}_{E,\Delta}$, it is preferable to use an ensemble $\rho_{B}$ with the same energy as $E$. In fig. 3(b), we selected the canonical ensemble with corresponding energies for different energy shells and calculated the average distinguishability as
$\sum_{E_{i}\in\mathcal{M}_{E,\Delta}}V_{\text{dg}}^{i}/d_{\mathcal{M}},$ (29)
where $d_{\mathcal{M}}$ denotes the number of energy levels within the energy shell. Note that this average distinguishability is not the same as the “diagonal” probabilistic typicality in eq. 19. The $\rho_{B}$ used in $V_{\text{dg}}^{i}$ may not be the microcanonical ensemble consisting of states within $\mathcal{M}_{E,\Delta}$. According to fig. 3(b), although the average distinguishability varies with the energy center $E$, the variation is significantly less compared to the changes caused by energy differences, as shown in fig. 3(a). Additionally, the average distinguishability in integrable systems remains higher than in non-integrable systems.
IV.2 The ETH under the local measurement
Figure 4: Parameters: $g=1.05$, $N_{B_{1}}=1$. (a) The average diagonal elements $\overline{V}^{\text{Loc}}_{\text{dg}}$ versus $N$. (b) The maximum diagonal element $V^{\text{Loc}}_{\text{dg},\max}$ versus $N$. (c) The average off-diagonal elements $\overline{V}^{\text{Loc}}_{\text{off}}$ versus $N$. (d) The maximum diagonal element $V^{\text{Loc}}_{\text{off},\max}$ versus $N$.
In this section, we focus on analyzing the strong and weak ETH under local measurements. We examine two energy shells $\mathcal{M}^{\prime}_{E=0,\Delta}$ and $\mathcal{M}_{E(\beta),\Delta}$, where $\Delta E=0.2N$ and $E(\beta)=E(\rho^{\text{c}}(\beta=0.1))$. Both energy shells include highly excited states. For the energy shell $\mathcal{M}$, we considered the canonical ensemble $\rho^{\text{c}}(\beta=0.1)$ and the mirocanonical ensemble $\rho^{\text{mc}}=\sum_{E_{i}\in\mathcal{M}}\Pi^{i}/d_{\mathcal{M}}$. For the energy shell $\mathcal{M}^{\prime}$, we used the mirocanonical ensemble ${\rho^{\prime}}^{\text{mc}}=\sum_{E_{i}\in\mathcal{M}^{\prime}}\Pi^{i}/d_{\mathcal{M}^{\prime}}$.
We compute the average distinguishability of the eigenstates within the energy shell as follows
$\overline{V}^{\text{Loc}}_{\text{(off-)dg}}=\frac{1}{d_{\mathcal{M}}}\sum_{i\in\mathcal{M}_{E,\Delta E}}V_{\text{(off-)dg}}^{i}.$ (30)
For two microcanonical ensembles, this average corresponds to the “(off-)diagonal” probabilistic typicality given in eq. 20 and eq. 19. To determine whether the diagonal elements $V_{\text{dg}}^{i}$ are exponentially small for all states in the energy shell, we calculate
$V^{\text{Loc}}_{\text{dg},\max}=\max_{E_{i}\in\mathcal{M}}V_{\text{dg}}^{i}.$ (31)
We also compute the maximum off-diagonal elements relative to the diagonal elements
$V^{\text{Loc}}_{\text{off},\max}=\max_{E_{i}\in\mathcal{M},j}V_{\text{off}}^{ij}.$ (32)
Here, as in eq. 13, the index $j$ spans the full Hilbert space.
According to eqs. 13 and 21, the local state of ensemble $\rho_{B_{1}}$ is a crucial quantity. As $N$ increases from $10$ to $15$, $\text{Tr}(\rho_{B_{1}}^{-1})-1$ shows almost no variation across several ensembles, regardless of whether the system is integrable. For $h=0$, we find $\text{Tr}((\rho_{B_{1}}^{\text{c}})^{-1})-1=3.04367676652\pm 10^{-11}$, $\text{Tr}((\rho_{B_{1}}^{\text{mc}})^{-1})-1=3.04\pm 10^{-2}$ and $\text{Tr}(({\rho^{\prime}}_{B_{1}}^{\text{mc}})^{-1})-1=3.0005\pm 5*10^{-4}$. For $h=0.1$, the values are $\text{Tr}((\rho_{B_{1}}^{\text{c}})^{-1})-1=3.04427703179\pm 10^{-11}$, $\text{Tr}((\rho_{B_{1}}^{\text{mc}})^{-1})-1=3.04\pm 10^{-2}$ and $\text{Tr}(({\rho^{\prime}}_{B_{1}}^{\text{mc}})^{-1})-1=3.0005\pm 5*10^{-4}$.
As shown in fig. 4(a-b), there is a significant difference between integrable and non-integrable systems in the behavior of the diagonal elements. In the non-integrable system, the average distinguishability of the diagonal elements decreases exponentially with $N$. The exponential decay constant is approximately $\frac{4}{5}\ln(2)$, which is closely aligns with the $\ln(2)=d\ln(1/D_{N})/dN$ derived in eq. 14. The exponential decay constant for the maximum $V_{\text{dg}}^{i}$ is slightly smaller, around $\frac{3}{5}\ln(2)$. In contrast, for the integrable system, the exponential decay constant for the average distinguishability is minimal, about $\frac{1}{5}\ln(2)$, and the maximum $V_{\text{dg}}^{i}$ seems to have almost no decay with $N$.
As shown in fig. 4(c-d), the difference in off-diagonal elements between integrable and non-integrable systems is small. Consistent with the prediction of eq. 14, the average off-diagonal elements are strictly suppressed exponentially, with an exponential decay constant of $\ln(2)$ in both cases, regardless of integrability.However, the exponential decay in the average off-diagonal elements does not imply that the maximum values of these elements will also decrease exponentially. As shown in fig. 4(d), although the maximum off-diagonal values do decrease with $N$, they do not exhibit a clear pattern.
Additionally, the average off-diagonal elements of the non-integrable system in fig. 4(c) are consistently smaller than those of the corresponding integrable system. This observation aligns with the trade-off relationship between the diagonal and off-diagonal elements described in eq. 21.
IV.3 The ETH under the average observable
Figure 5: Parameters: $g=1.05$, $N_{B_{1}}=1$. (a) The average diagonal elements $\overline{V}^{\text{Avg}}_{\text{dg}}$ versus $N$. (b) The maximum diagonal element $V^{\text{Avg}}_{\text{dg},\max}$ versus $N$. (c) The average off-diagonal elements $\overline{V}^{\text{Avg}}_{\text{off}}$ versus $N$. (d) The maximum diagonal element $V^{\text{Avg}}_{\text{off},\max}$ versus $N$.
Now, let us explore the strong and weak ETH under the average observable. The energy shells and ensembles are consistent with those in the previous section. Given that the ensembles used here are all translation-invariant, we simply replace $V(\rho_{B},O^{B_{1}}_{ij})$ with $V(\rho_{B},\overline{O}^{B_{1}}_{ij})$ to derive the corresponding $\overline{V}^{\text{Avg}}_{\text{(off-)dg}}$ and $V^{\text{Avg}}_{\text{(off-)dg},\max}$ from eqs. 30, 31 and 32. According to sections III.3 and 25, the local term is obtained by multiplying $\text{Tr}(\rho_{B_{1}}^{-1})-1$ from the previous section by a factor of $1/\mathcal{C}$. In the case of the average observable, the spatial correlation of the ensemble states, represented by $V^{\text{Avg}}_{\text{Cor}}$, also needs to be considered. As shown in fig. 2, the correlation term is negative, which further lowers the upper bound of the average (off-)diagonal elements. Additionally, the correlation terms vary significantly depending on the energy center, as the ensemble $\rho^{\text{mc}}$ with energy near zero exhibits weaker correlations. This difference also accounts for the more pronounced distinction between the two ensembles in fig. 5(c) compared to fig. 4(c). Although the localized states of the ensembles are quite similar, their correlation terms differ.
As shown in fig. 5(a-d), the significant difference in the exponential decay constants of the diagonal elements between integrable and non-integrable systems remains evident. The average of the off-diagonal elements continues to be exponentially suppressed, and the maximum values of the off-diagonal elements still display no clear pattern. However, in fig. 5(c), the average off-diagonal elements of the non-integrable system are no longer consistently smaller than those of the corresponding integrable system. This observation is still consistent with the trade-off relationship between the diagonal and off-diagonal elements given by eq. 25, as fig. 5(a) shows that the average diagonal elements of the integrable system are now not always larger than those of the non-integrable system.
V Conclusion and discussion
In this paper, we delve deeper into the observable-independent and versatile distinguishability measure $V(\rho_{B},O^{B_{1}}_{ij})$. We explore how this measure can be integrated into the framework of subsystem ETH and weak ETH, and extend it to apply to average observables. By modifying this measure, we identify a precise and universal trade-off relationship between its diagonal and off-diagonal elements. Numerical calculations confirm that this trade-off relationship also holds for other ETH measures, albeit not as an equality. The trade-off applies not only to common ensembles but also to a wide range of states $\rho_{B}$. The right side of the trade-off equation depends solely on the local state of $\rho_{B}$, while the left side includes a diagonal element and $D_{N}-1$ off-diagonal elements, enforcing exponential suppression of the average off-diagonal elements. The diagonal element’s significance lies in its ratio to the average off-diagonal elements, which relates to the system’s integrability. We also establish the trade-off relationship for weak ETH and average observables between their respective diagonal and off-diagonal elements. Our numerical results demonstrate that the measure $V(\rho_{B},O^{B_{1}}_{ij})$ is indeed effective for analyzing ETH, and the strict exponential suppression of the average off-diagonal element aligns with the trade-off relationship’s predictions.
In the numerical calculations presented in this paper, we compare eigenstates within an energy shell using the same ensemble. However, the energies of these states may deviate from that of the ensemble, introducing additional distinguishability, as seen in fig. 3(a). By using an ensemble with an energy that matches the energy eigenstates, we can further reduce both the average and maximum distinguishability. Additionally, examining how the energy width of the microcanonical ensemble and interactions influence the ratio between diagonal and average off-diagonal elements is crucial for deepening our understanding of the theoretical mechanisms behind the ETH.
Additionally, this paper does not delve into cases where the size of subsystem $B_{1}$ scales proportionally with the full system, $N_{B_{1}}/N_{B}=p$, as this has already been thoroughly discussed in [14]. However, using the measure form $d^{(3)}$ from this paper, there are still aspects worth exploring. For local observations, since $\text{Tr}(\rho_{B_{1}}^{-1})-1$ generally increases with $D_{B_{1}}^{2}$, if the ratio of diagonal to average off-diagonal elements $f^{i}(N)$ increases exponentially at a much slower rate than $D_{N}$, then we would have $V_{\text{dg}}^{i}\propto D_{B_{1}}^{2}/D_{N}$. According to section II.1, we obtain $d^{(2)}\sim D_{B_{1}}/D_{N}^{1/2}$, which aligns with the results in [14]. Similarly to appendix A, we find that
$V(\rho^{\text{c}}_{B},O^{B_{1}}_{ii})=\sum_{\alpha,\beta}{p^{\prime}_{\beta}}^{-1}\lvert\bra{\beta}(\sigma^{B_{1}}_{i}-\rho_{B_{1}}^{\text{c}})\ket{\alpha}\rvert^{2}$
$\leq\sum_{\alpha,\beta}{p^{\prime}_{\beta}}^{-1}(\lvert\bra{\beta}(\sigma^{B_{1}}_{i}-\rho_{B_{1}}^{\text{mc}})\ket{\alpha}\rvert+\lvert\bra{\beta}(\rho_{B_{1}}^{\text{mc}}-\rho_{B_{1}}^{\text{c}})\ket{\alpha}\rvert)^{2}.$ (33)
When $B_{1}$ is scaled to infinity, a semiclassical analysis similar to that in [14] can be performed to approximate $V(\rho^{\text{c}}_{B},O^{B_{1}}_{ii})$. This provides an alternative perspective for considering the local differences between energy eigenstates and the canonical ensemble in the thermodynamic limit.
Acknowledgements.
We would like to thank Xiao-Kan Guo for his helpful comments. This work is supported by the National Natural Science Foundation of China under Grant No. 12305035.
Appendix A Variance relation
Here, we present the detailed derivation of the key quantum variance relation given in eqs. 13 and III.3 of the main text. Starting with the definition
$V(\rho_{B},O^{B_{1}}_{ij})=\lVert(\sigma^{ij}_{B_{1}}-\delta_{i,j}\rho_{B_{1}})\rho^{-1/2}_{B_{1}}\rVert^{2}_{2}$
$=\text{Tr}(\sigma^{ij}_{B_{1}}\rho^{-1}_{B_{1}}\sigma^{ij{\dagger}}_{B_{1}})-\delta_{i,j}.$ (34)
Assuming the spectral decomposition $\rho_{B_{1}}=\sum_{\alpha}p^{\prime}_{\alpha}\Pi^{\alpha}_{B_{1}}$, we can express
$V_{\text{dg}}^{i}+\sum_{j;j\neq i}V_{\text{off}}^{ij}=\sum_{j,\alpha,\beta}{p^{\prime}_{\beta}}^{-1}\lvert\text{Tr}(\sigma^{ij}_{B_{1}}\Pi^{\alpha\beta}_{B_{1}})\rvert^{2}-1,$ (35)
where $\Pi^{\alpha\beta}=\ket{\alpha}\bra{\beta}$ |. Recognizing that $\lvert\text{Tr}(\sigma^{ij}_{B_{1}}A^{B_{1}})\rvert^{2}=\lvert\text{Tr}(\Pi^{ij}A^{B_{1}})\rvert^{2}$, and using
$\mathcal{J}_{\rho_{B_{k}}}^{-1/2}(\Pi^{\alpha\beta}_{B_{k}})=(p^{\prime}_{\alpha}p^{\prime}_{\beta})^{-1/2}\Pi^{\alpha\beta}_{B_{k}}$ (36)
along with the definition
$O^{B_{1}}_{\alpha\beta}=\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\Pi^{\alpha\beta}_{B_{1}}),$ (37)
we can further rewrite
$V_{\text{dg}}^{i}+\sum_{j;j\neq i}V_{\text{off}}^{ij}=\sum_{j,\alpha,\beta}{p^{\prime}_{\beta}}^{-1}\lvert\text{Tr}(\Pi^{\alpha\beta}_{B_{1}}\Pi^{ij}_{B})\rvert^{2}-1$
$=\sum_{\beta,\alpha}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{1}}_{\alpha\beta})^{\dagger}\Pi_{i}O^{B_{1}}_{\alpha\beta}]-1$
$=\sum_{\beta}{p^{\prime}_{\beta}}^{-1}\text{Tr}(\rho_{B_{1}}O^{B_{1}}_{ii})-1=\text{Tr}(\rho_{B_{1}}^{-1})-1,$ (38)
which directly corresponds to the relation in eq. 13.
The proof of section III.3 follows similar approach as appendix A. The key difference is the substitution of $O$ with $\overline{O}$, and the use of the relation
$\lvert\text{Tr}[(\frac{1}{\mathcal{C}}\sum_{k}\sigma^{ij,k}_{B_{1}})A^{B_{1}}]\rvert^{2}=\lvert\text{Tr}[\Pi^{ij}(\frac{1}{\mathcal{C}}\sum_{k}A^{B_{k}})]\rvert^{2}.$ (39)
Here, we need to use $\overline{O}^{B}_{\alpha\beta}=\frac{1}{\mathcal{C}}\sum_{k}O^{B_{k}}_{\alpha\beta,1}$, where $O^{B_{k}}_{\alpha\beta,1}$ represents the translational copies of $O^{B_{1}}_{\alpha\beta}$. Using this, we can rewrite the expression as
$V(\rho,\overline{O}^{B_{1}}_{ii})+\sum_{j;j\neq i}V(\rho,\overline{O}^{B_{1}}_{ij})$
$=\sum_{\beta,\alpha}p^{\prime}_{\alpha}\text{Tr}[(\overline{O}^{B}_{\alpha\beta})^{\dagger}\Pi_{i}\overline{O}^{B}_{\alpha\beta}]-1$
$=\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,k}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\Pi_{i}O^{B_{k}}_{\alpha\beta,1}]-\frac{1}{\mathcal{C}}$
$+\frac{1}{\mathcal{C}^{2}}\sum_{\beta,\alpha,l,k;k\neq l}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\Pi_{i}O^{B_{l}}_{\alpha\beta,1}]-\frac{\mathcal{C}-1}{\mathcal{C}}.$ (40)
In this expression, the local terms are given by
$\sum_{\beta,\alpha}p^{\prime}_{\alpha}\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\Pi_{i}O^{B_{k}}_{\alpha\beta,1}]$
$=\sum_{\beta}{p^{\prime}_{\beta}}^{-1}\text{Tr}(\rho_{B_{1}}\mathcal{J}_{\rho_{B_{1}}}^{-1/2}(\sigma^{ii,k}_{B_{1}}))=\text{Tr}(\rho_{B_{1}}^{-1}).$ (41)
Combining this with
$\text{Tr}[(O^{B_{k}}_{\alpha\beta,1})^{\dagger}\sigma^{ii}_{B_{k}}]\times\text{Tr}(\rho^{1}_{B_{l}}O^{B_{l}}_{\alpha\beta,1})={p^{\prime}_{\alpha}}^{-1}\text{Tr}[\Pi_{\alpha}\sigma^{ii}_{B_{k}}]\delta_{\alpha\beta},$ (42)
we arrive at section III.3.
Figure 6: Parameters: $g=1.05$, $N=15$. (a) The number of states of the spin chain system. (b) Energy probability distribution of the canonical ensemble with $\beta=0.1$. (c) Energy as a function of the thermodynamic beta. (d) The value of $\log_{10}I(B_{1},B_{2})$ between two sites at $\beta=0.1$.
Appendix B Basic properties of the model
Here, we briefly illustrate the basic properties of the model. As shown in fig. 6(a), the density of states for both integrable and non-integrable systems is quite similar, with the non-integrable system exhibiting a smoother distribution that can be reasonably approximated by a binomial distribution [14]. The energy-temperature curve for the system is provided in fig. 6(c). Additionally, as shown in fig. 6(d), the mutual information $I(B_{1},B_{2})$ decays exponentially with the distance between sites. The plot depicts the system with $h=0.1$; the mutual information for the $h=0$ system is nearly identical and therefore not shown.